OpenStudy (anonymous):

Evaluate the expression below when x = -2 X^2-8/x+6 Can anyone help me (:

OpenStudy (anonymous):

So you'll start with (-2)^2 - 8 to find the top, then you'll do (-2) + 6 to find the bottom. After that, simply divide the top you found by the bottom you found

OpenStudy (igreen):

Input 2 for every x: \(\dfrac{x^2 - 8}{x+6}\) \(\dfrac{(2)^2 - 8}{2+6}\) Simplify the exponent: \(\dfrac{4 - 8}{2+6}\) Can you do the rest? @JulianoRossiel

OpenStudy (jebonna):

I don't know much about algebra but I know.. 1+1=2...